OpenStudy (anonymous):

Jen buys a motorcycle for $11,000. The dealer is charging her an annual interest rate of 10.75% and is using the add-on method to compute monthly payments. If she pays off the loan in 84 months, what are her monthly payments? If she makes a down payment of $2,000, how much will her monthly payments be? Do not round until the final answer. Then, round to the nearest cent. A $459.90; $187.77 B $229.49; $187.77 C $229.49; $376.29 D $229.49; $271.76

OpenStudy (campbell_st):

well 84 months = 7 years and the interest is per annum with a $2000 deposit i means there is $9000 to be repaid find the simple interest as the 1st step I = 9000*10.75/100* 7 then the interest is I = 6772.50 total to be repaid is $9000 + 6772.50 = 15772.50 monthly repayment are 15772.50/84 = $187.77

OpenStudy (anonymous):

Thank you again, very helpful!